News reports speak of an emerging crisis of childhood obesity in the United States. The National Health and Nutrition Examination Survey (NHANES) is a government survey run every several years recording a number of vital statistics on a random sample of Americans. A body mass index (BMI) is computed for each individual in the sample based on the individual's height and weight. Here are sample results for 8-year-old boys over the past 40 years. The table gives the sample size for 3 different surveys, the sample mean and sample standard deviation. Suppose we want to run an ANOVA to see if there is a difference in mean BMI for the three surveys. Assume each survey is a random sample and there are no extreme outliers.
Answer #1
Concepts and reason

Analysis Of Variance (ANOVA) is a statistical model used to analyse the difference between the group means. It is a way to find the significant results in an experiment. It is used when there are more than two groups.

The associated procedures developed by statistician and evolutionary biologist Ronald Fisher.

Fundamentals

The overall mean anx+nx, un,
n+ n +n,

SSG =n, (, - x) +n (7z - ) +n (-x)?

The formula for the error sum of squares is as follows:

SSE = SST - SSG

The formula of the degrees of freedom for treatment sum of squares is as follows:

df, =k-1

The formula of the degrees of freedom for error sum of squares is as follows:

df.=n-k

The formula of degrees of freedom for total sum of squares is as follows:

ofy =n-1

The formula of the mean sum of squares for the treatments is as follows:

MSG =
SSG
dfg

The formula for the mean sum of squares for the error is as follows:

The formula of the F statistic is as follows:

MSE
F=
MSG

(a)

The sample size is large greater than 30. So, use ANOVA technique to see if there is any difference in mean BMI for the three surveys.

(b)

The given information is as follows:

Survey
NHES II, 1965 618 16.3 2.486
NHANES II 1980 145 16.5 2.408
NHANES II 2002 214 18.4 5.851

Therefore, the sample standard deviations of each survey are:

1.The NHES II 1965 is 2.486

2.The NHANES II 1980 is 2.408

3.The NHANES II 2002 is 5.851

(c)

Null hypothesis, all group means are equal.

Alternative hypothesis, all group means are not equal.

The overall mean r-ny+nx, +nex
n+h+₂
16403.5
977
=16.7897

The sum of squares for groups
SSG = n, (6, – x) ° + n3 (52 – x)* +n, (75 – x)
= 618(16.3-16.7897) + +214(18.4-16.7897)
= 148

The degrees of freedom for groups = 3-1
= 2

Therefore, the mean sum of squares for groups,

SSG
18907.74
= 9453.87

The sum of squares for error is,

SSE= (n -1).s+(n2-1)sz? +(nz-1).s?
(618-1)x2.486 + +(1079–1)x5.851
= 3813.181+834.9788+7291.885
= 11940.04

Degrees of freedom for error is,

= 977-3
| = 974

The mean sum of squares for error is,

SSE -11940.04
974
= 12.2588

The test statistic is,

MSG
F-ratio =
MSE
9453.87
12.2588
=771.1905

[Part c]

Ans: Part a

The sample size is large .There are no extreme outliers in the given information. Hence, use ANOVA technique to see if there is any difference in mean BMI for the three surveys.

Part b

The sample standard deviations of each survey are:

1.The NHES II 1965 is 2.486

2.The NHANES II 1980 is 2.408

3.The NHANES II 2002 is 5.851

Part c

The ANOVA table is as follows:

Mean sum
Sum of
as Mean sum
F
Source of Sum of
Variation Squares
Groups 9453.872
Error 11940.04 974
Total 21393.91 976
771.19

From Excel=FDIST(771.1905,2,974)
, the value of p is 0.00000

Compare the p-value with the level of significancea=0.05
.

Here, the p-value is lesser than the level of significance then reject the Null hypothesis.

Use p-value to conclude the decision.

Hence, conclude that all treatment means are not equal.